OpenStudy (anonymous):

A country's population, in millions, is given by the exponential model P = 78.2e0.047t, where t represents years after 2000. What population does the model predict for the country in 2012? (Round answer to the nearest tenth.)

OpenStudy (anonymous):

f t represents years after 2000 and you would like to know the population in the year 2012, first calculate the t value, in this case: 2012-2000 = 12 This is the only unknown number in your formula, as e is a mathematical constant that approximates 2,718. So you formula then looks like this: P = 78,2 * 2,718 ^ (0,047 * 12) then: 0,047 * 12 = 0,564 which gives: P=78,2 * 2,718 ^ 0,564 First the power function: 2,718 ^ 0,564 = 1,758...... which you then multiply with 78,2 P = 78,2 * 1,758.... Which eventually gives a population of: P = 137,451...... Rounding it to the nearest tenth will give P= 137,5